OpenStudy (anonymous):

Solve the following using quadratic equations: x^2+ 6x + 3 = 0

OpenStudy (anonymous):

Would you like the quadratic formula or just the answer?

OpenStudy (anonymous):

Just answer'

OpenStudy (anonymous):

x = sqrt(6) - 3, x = -3-sqrt(6)

OpenStudy (anonymous):

x^(2)+6x+3=0 Use the quadratic formula to find the solutions. In this case, the values are a=1, b=6, and c=3. x=(-b\~(b^(2)-4ac))/(2a) where ax^(2)+bx+c=0 Substitute in the values of a=1, b=6, and c=3. x=(-6\~((6)^(2)-4(1)(3)))/(2(1)) Simplify the section inside the radical. x=(-6\2~(6))/(2(1)) Simplify the denominator of the quadratic formula. x=(-6\2~(6))/(2) First, solve the + portion of \. x=(-6+2~(6))/(2) Simplify the expression to solve for the + portion of the \. x=-3+~(6) Next, solve the - portion of \. x=(-6-2~(6))/(2) Simplify the expression to solve for the - portion of the \. x=-3-~(6) The final answer is the combination of both solutions. \[x=-3+\sqrt6,-3-\sqrt6\] ~ means sqrt